### v1

# DRP: Distilled Reasoning Pruning with Skill-aware Step Decomposition for Efficient Large Reasoning Models

Abstract
While Large Reasoning Models (LRMs) have demonstrated remarkable success in complex reasoning tasks through Long Chain-of-Thought (CoT) reasoning, their inference often involves excessively verbose reasoning traces, resulting in substantial inefficiency. To address this issue, we propose Distilled Reasoning Pruning (DRP), a hybrid framework that combines inference-time pruning with tuning-based distillation—two widely used strategies for efficient reasoning. DRP employs a teacher model to perform skill-aware step decomposition and content pruning, and then distills the pruned reasoning paths into a student model, enabling it to reason both efficiently and accurately. Across a series of challenging mathematical reasoning datasets, we find models trained with DRP achieve substantial improvements in token efficiency without sacrificing accuracy. Specifically, DRP reduces the average token usage on GSM8K from 917 to 328 while improving accuracy from 91.7% to 94.1%, and achieves a 43% token reduction on AIME with no performance drop. Further analysis reveals that aligning the reasoning structure of training CoTs with the student’s reasoning capacity is critical for effective knowledge transfer and performance gains.
DRP: Distilled Reasoning Pruning with Skill-aware Step Decomposition for Efficient Large Reasoning Models
Yuxuan Jiang1  Dawei Li2  Francis Ferraro1 1University of Maryland, Baltimore County 2Arizona State University yuxuanj1@umbc.edu
1 Introduction
Although Large Reasoning Models (LRMs) Xu et al. (2025a), like OpenAI’s o1 OpenAI (2024b) and DeepSeek-R1 Guo et al. (2025), have advanced the state of the art in complex reasoning tasks Li et al. (2025b), a critical limitation of these models is their tendency toward overthinking—the generation of excessively verbose reasoning trajectories containing redundant or unnecessary steps Chen et al. (2024); Cui et al. (2025); Fu et al. (2024). This can lead to substantial inference overhead and misguide the model toward incorrect conclusions Sui et al. (2025).
Existing solutions predominantly follow two paradigms: inference-time pruning, which attempts to terminate generation early to avoid redundant reasoning steps Cui et al. (2025); Muennighoff et al. (2025); Fu et al. (2024); Zeng et al. (2025), and distillation-based compression, where smaller models are trained on teacher-generated reasoning paths to imitate the concise reasoning behavior of larger models Xu et al. (2025b); Zhu et al. (2024). However, both approaches can be at the cost of accuracy: pruning methods risk prematurely halting the reasoning process, thereby preventing the model from reaching the correct answer, while distillation methods tend to underperform—particularly for small-scale student models—due to the learnability gap, which hinders effective knowledge transfer when the performance or size gap between student and teacher is too great Xu et al. (2024); Li et al. (2025a).
Figure 1: An overview of our proposed Distilled Reasoning Pruning framework, which unifies pruning and distillation by having the teacher model prune the student model’s reasoning chains and distill the resulting concise CoT back into the student. This approach improves reasoning efficiency without sacrificing accuracy.
Figure 2: Overview of our proposed DRP framework, which distills concise, skill-aware reasoning paths into a target model to improve reasoning efficiency while preserving accuracy.
To address these limitations, we propose Distilled Reasoning Pruning (DRP) (Fig. 1), a hybrid framework that combines the strengths of both pruning and distillation (student/teacher) paradigms. In particular, rather than simply distilling down a response from the teacher model, we use the teacher model to prune an initial, lengthy, CoT reasoning trajectory from the student model. To facilitate this, and to encourage shorter yet informative resulting pruned trajectories, we introduce a skill-based step decomposition method. The teacher model uses this to prune the trajectory, which produces more stable and semantically coherent reasoning units. Unlike prior methods that rely solely on either teacher-generated Xu et al. (2025b); Zhu et al. (2024) or self-sampled Chen et al. (2024); Ma et al. (2025) concise trajectories for distillation, DRP takes advantage of the teacher’s pruning within the student model’s original reasoning structure. This design reduces the learnability gap and enables student models to achieve efficient reasoning without compromising performance.
Consider the example in Fig. 2. For this word problem, the student model generates its initial CoT trajectory (shown in the <think> block). This trajectory is provided to a teacher model, which segments that CoT into steps, with a high-level description of the skill that that step demonstrates or exercises. The teacher model then prunes and curates these skill-segmented steps, such as by merging similar ones (if the student was redundant or verbose) or deleting steps (e.g., for backtracking). This pruned CoT trajectory is then provided to the student model for supervised fine-tuning.
We conduct extensive experiment in various student models, teacher models and mathematical datasets. DRP consistently improves the efficiency and accuracy of student models—even on significantly harder out-of-distribution tasks. Our ablation study demonstrates that DRP substantially outperforms direct distillation approaches. In addition, we show that our framework consistently benefits from various teacher models. We believe all these findings and insights from DRP can benefit future works in efficient reasoning in LRMs. Our key contributions are as follows:
  • We propose Distilled Reasoning Pruning, a novel framework that unifies step-level pruning and distillation to improve both reasoning efficiency and accuracy in small-scale LLMs.
  • We introduce a skill-based step decomposition method that segments reasoning traces into semantically coherent and functionally aligned units, providing a stable foundation for pruning and supervision.
  • We point out that effective training CoTs should be both informative and structurally consistent with the student’s reasoning process, thereby facilitating knowledge transfer and bridging the capability gap between teacher and student models.
2 Related Work
2.1 The Overthinking Problem
Large reasoning models, such as OpenAI o1 OpenAI (2024b) and DeepSeek-R1 Guo et al. (2025), are a specialized subclass of large language models trained to iteratively generate intermediate reasoning steps and refine the thinking process until reaching a final answer Chen et al. (2025); Sui et al. (2025). Unlike traditional LLMs, which rely on test-time strategies such as Chain-of-Thought prompting to elicit detailed reasoning, LRMs are trained to internalize CoT generation. This design significantly enhances their performance on reasoning-intensive tasks, particularly in challenging mathematics and programming. However, LRMs, especially those with smaller parameter sizes, often exhibit a tendency to produce overly verbose reasoning chains, including unnecessary tracebacks and redundant alternative paths Chen et al. (2024); Sui et al. (2025); Cui et al. (2025); Fu et al. (2025, 2024). Such overthinking not only increases token consumption beyond practical inference budgets but also introduces reasoning drift that can degrade output correctness Hou et al. (2025); Sui et al. (2025). To address these, our method explicitly prunes harmful overthinking behaviors while simultaneously guiding the model toward concise and effective reasoning expressions.
2.2 Token-Efficient Reasoning Methods
Current token-efficient reasoning methods can be categorized into three approaches Sui et al. (2025):
(1) Prompt-based methodsimpose token budget constraints within the prompt to encourage concise reasoning without altering model parameters. For example, TALE Han et al. (2024) estimates per-instance token budgets to reduce length while maintaining accuracy on simple tasks. However, these methods rely on handcrafted prompt designs and struggle with complex inputs. (2) Supervised fine-tuning (SFT)approaches train models on compressed reasoning traces to promote internalized efficiency Xia et al. (2025); Munkhbat et al. (2025). CoT-Valve Ma et al. (2025), for instance, uses iterative compression stages to fine-tune reasoning conciseness. Yet, these approaches often require task-specific data and retraining. (3) Reinforcement learning (RL)methods guide models toward concise reasoning by introducing explicit reward signals that penalize overly long outputs Team et al. (2025); Chen et al. (2024). Some approaches incorporate early-exit mechanisms at inference time Muennighoff et al. (2025); Fu et al. (2024); Zeng et al. (2025). A representative method is ThinkPrune Hou et al. (2025), which defines a target reasoning length and progressively fine-tune models under increasingly tighter constraints.
Despite progress, existing approaches often trade off accuracy, especially in out-of-domain (OOD) tasks, for brevity. In contrast, we propose a lightweight framework that leverages external teacher models to perform skill-aware pruning over reasoning steps. This structure-aware intervention enables significant token reduction while improving performance, especially under distribution shift.
3 Methodology
We propose Distilled Reasoning Pruning (DRP), a method that distills concise reasoning paths into a student model by first decomposing them into skill-based steps and then pruning redundant content with the help of a teacher model (e.g., GPT-4o). As illustrated in Figure 2, DRP proceeds in three stages: (1) decomposing the original reasoning into fine-grained, skill-based steps; (2) pruning and rewriting these steps via a teacher LLM; and (3) fine-tuning the student model on the revised, compact trajectories. These stages result in token-efficient supervision that improves both reasoning efficiency and accuracy during fine-tuning.
3.1 Skill-Based Step Decomposition
To solve a math problem, a reasoning model generates a token sequence as the full response, which consists of two components: a structured thinking process $T$, enclosed within the <think> </think> tags, and a final answer summarization $A$, captures the conclusive statement derived from the reasoning trace. Together, we terms these a “response,” $R=(T,A)$. For instance, in example shown in Fig. 2, the “thinking” portion of the initial response would include “Natalia sold 48 clips in …Total clips sold: 48 + 24 = 72” while the answer summarization is “The answer is 72.”
We extract the thinking segment $T$ from the student model’s response and prompt a teacher model to decompose it into a sequence of non-overlapping segments, each aligned with a functional reasoning skill such as arithmetic computation, comparison, or logical inference. This skill-based decomposition serves as the foundation for step-level pruning and distillation. Compared to naïve sentence-based splitting, it enhances the stability of step boundaries, maintains semantic coherence across reasoning steps, and provides a more fine-grained supervision signal (in § 6.1). We validate the quality of this segmentation through a pairwise evaluation using Gemini 2.0 Flash as an external judge; results and examples are presented in Appendix D for review.
Formally, given a response $R=(T,A)$, we define a decomposition function $D$ that maps the thinking trace into skill-annotated steps:
$D(T)\mapsto\{(s_{1},k_{1}),(s_{2},k_{2}),\ldots,(s_{m},k_{m})\},$
where each $s_{i}$ denotes a contiguous span of tokens representing a single reasoning step, and $k_{i}$ is the associated skill label, inferred by the teacher model based on its internal knowledge and understanding of the reasoning process. For the example in Fig. 2, one segment should be “Natalia sold 48 clips in April,” which corresponding to the skill “Reading given quantity.” However, these skills can cover broader topics such as “Algebraic representation” and “Interpret a fraction of a subset.” The complete decomposition prompts are provided in Appendix C.
3.2 Step-Level Pruning
For each reasoning step–skill pair $(s_{i},k_{i})$, the teacher model is prompted to revise the step without altering the essential structure or logical intent of the original reasoning. The teacher selects one of the following actions for each step:
  • Keep: Retain the step unchanged.
  • Delete: Remove the step if it is redundant or uninformative.
  • Rewrite: Replace the step with a more concise version conveying the same logic.
  • Merge: Combine the step with adjacent ones if they form a coherent atomic unit.
This yields a revised step $\hat{s}_{i}=\text{Revise}(s_{i})$, and a pruned reasoning trace: $\hat{T}=\{\hat{s}_{1},\hat{s}_{2},\dots,\hat{s}_{m^{\prime}}\}$, where $m^{\prime}\leq m$, which reduces redundancy and increases the overall information density.
Finally, the teacher model rewrites $\hat{T}$ into a fluent, coherent reasoning trace that preserves the tone and speaking style of the student model. To ensure consistency between the revised reasoning and the final answer, we prompt the teacher to optionally revise the original answer segment $A$, yielding an updated final answer summarization $\hat{A}$. The final output becomes a concatenation $\hat{R}=(\hat{T},\hat{A})$, which we use as the target for supervised fine-tuning. Complete prompt are in Appendix C.
Self-Revision Prompt Given a list of reasoning steps labeled with their respective skills, your task is to evaluate and revise each step according to one of the four ACTIONs. Ensure the resulting reasoning path is concise, fluent, and logically sound. At the end, synthesize the revised steps into a coherent explanation that matches the speaker’s tone.
3.3 Supervised Fine-Tuning
We construct the training dataset using pairs $(x,\hat{R})$, where $x$ is the input question and $\hat{R}$ is the complete revised response.
We fine-tune the model using teacher-forced decoding to encourage the generation of concise reasoning traces. The training objective minimizes the negative log-likelihood of the revised response:
$\mathcal{L}_{\text{SFT}}=-\sum_{i=1}^{n}\log P_{\theta}(y_{i}\mid x,y_{<i}),$
where $\{y_{1},\dots,y_{n}\}$ are the tokens in $\hat{R}$, and $\theta$ denotes the model parameters. This supervision enables the model to internalize skill-aligned, token-efficient reasoning strategies while maintaining consistency with the final answer.
Method GSM8K MATH500 (OOD) AIME24 (OOD) AMC (OOD)
Pass@1 #Tokens Pass@1 #Tokens Pass@1 #Tokens Pass@1 #Tokens
R1-Distill-Qwen-1.5B
Base 70.7% 1443 80.4% 3276 6/30 10484 23/40 6516
+TALE 70.1% 1170 76.2% 3107 6/30 8915 22/40 6235
+Cot Valve 70.4% 805 76.5% 2705 7/30 5601 22/40 4574
+Thinkprune 80.0% 712 79.2% 2006 9/30 5745 25/40 3291
+DRP 83.4% 721 (-50%) 82.0% 2122 (-35%) 10/30 6135 (-42%) 27/40 3657 (-44%)
R1-Distill-Qwen-7B
Base 91.7% 917 92.4% 2486 15/30 8674 31/40 4845
+TALE 91.0% 522 91.6% 2530 10/30 8602 31/40 3998
+Cot Valve 90.8% 364 89.4% 1975 13/30 6315 30/40 3157
+DRP 94.1% 328 (-64%) 93.0% 1781 (-28%) 15/30 4966 (-43%) 33/40 3258 (-33%)
Table 1: Pass@1 accuracy and average token usage on R1-Distill-Qwen models across various math benchmarks, comparing our DRP method with Cot Valve Ma et al. (2025), TALE Han et al. (2024), and ThinkPrune Hou et al. (2025).
4 Experimental Setup
4.1 Datasets
Our training corpora consist of the training split of GSM8K Cobbe et al. (2021) and the full PRM12K Lightman et al. (2023) dataset. From these, we generate initial reasoning paths, which are then processed through skill-based step decomposition and teacher-guided step-level pruning to create supervision signals for fine-tuning. To evaluate complex mathematical reasoning and generalization ability, we select a broad set of out-of-domain benchmarks including MATH500 Hendrycks et al. (2021), AIME24 AI-MO Team (2024a), and AMC23 AI-MO Team (2024b).
4.2 Models
We use DeepSeek-R1-Distill-Qwen-7B and DeepSeek-R1-Distill-Qwen-1.5B Guo et al. (2025) as student models for supervised fine-tuning. Both are distilled variants of DeepSeek-R1 optimized for efficient inference. Our primary teacher model is GPT-4o OpenAI (2024a), which performs step-level decomposition and pruning to generate supervision signals. For ablation studies, we additionally explore alternative teacher models, including Gemini 2.0 Flash DeepMind (2025) and ChatGPT OpenAI (2022).
4.3 Compared Methods
We select three representative methods that span the major paradigms for token-efficient reasoning: (1) TALEHan et al. (2024): a prompt-based method incorporates a soft token budget constraint into the prompt to encourage concise generation. (2) CoT-ValveMa et al. (2025): a SFT-based method which enerates multiple chains-of-thought of varying lengths for the same problem, and performs supervised fine-tuning in multiple rounds—each time using shorter CoTs as training targets. (3) ThinkPruneHou et al. (2025): a tuning-based method uses reinforcement learning to iteratively reduce chain-of-thought length by optimizing under a target token constraint.
4.4 Implementation Details
Supervised fine-tuning is performed using the LLaMA-Factory [^1] https://github.com/hiyouga/LLaMA-Factory framework with LoRA adaptation. All models are trained for 3 epochs with cosine learning rate scheduling. Full hyperparameters are listed in Appendix B.
4.5 Evaluation Protocol
We use the lmms-eval [^2] https://github.com/EvolvingLMMs-Lab/lmms-eval framework for unified evaluation across tasks. Each model is evaluated in a zero-shot setting. For inference, we use the vLLM backend with the maximum generation length set to 131,072 tokens—the upper limit supported by the Qwen models.
4.6 Evaluation Metrics
We report Pass@1 as the accuracy metric. For efficiency, we measure the number of reasoning tokens generated per completion using the HuggingFace-compatible Qwen tokenizer. [^3] https://huggingface.co/Qwen/Qwen-tokenizer
We also observe that models occasionally fall into degenerate loops, repeatedly generating parts of their responses until reaching the maximum generation limit (e.g., 130k tokens), far exceeding the typical average length (e.g., 5k). In most of these cases, the model fails to answer correctly. Such outliers significantly inflate the average token count.
To mitigate their impact, we set a cutoff threshold of 12k tokens, which empirically covers 99% of correct responses across all benchmarks for the models we evaluate. Detailed token length distributions by task are provided in Appendix A.
5 Main Results
5.1 Accuracy and Token Efficiency
Table 1 presents the main results on our DRP and other compared methods. Our key findings are:
DRP consistently reduces token usage across all benchmarks and model sizes.
Our proposed DRP method achieves substantial reductions in average token usage on both in-domain and out-of-domain tasks. On GSM8K, DRP reduces token count by up to 64% with the 7B model. For out-of-domain datasets, DRP yields 28%–44% reductions across all benchmarks, demonstrating strong generalization beyond the training distribution.
DRP improve the accuracy by mitigating the over-thinking problem in LRMs.
Despite significantly reducing token usage, DRP preserves or improves Pass@1 accuracy on nearly all benchmarks. Notably, DRP improves accuracy even on harder datasets such as AMC and MATH500. The only exception is AIME24 under the 7B setting, where accuracy remains unchanged, suggesting the inherent difficulty of this benchmark.
Accuracy gains are more pronounced for smaller models.
DRP delivers particularly strong improvements with the 1.5B model. On GSM8K, accuracy increases by 12.7%, while on AIME24 and AMC, DRP answers 4 more problems correctly compared to the base model. This indicates DRP’s effectiveness in compensating for limited model capacity through more efficient supervision.
5.2 Comparison to Prior Work
We compare our approach with three representative baselines covering prompt-based, SFT-based, and RL-based token-efficient reasoning strategies (details can be found in Section 4.3).
Prompt-based methods offer limited control on Complex Reasoning Tasks.
TALE is simple to implement and demonstrates moderate effectiveness on short-answer tasks such as GSM8K. It achieves token reductions with minimal accuracy degradation (less than 1%) on both model sizes. However, its effectiveness diminishes significantly on more challenging benchmarks. For example, on AIME24 with the 7B model, TALE causes a notable performance drop. Overall, TALE’s prompt constraints provide limited control over fine-grained reasoning behaviors, which we hypothesize leads to its suboptimal performance on tasks requiring deeper or longer reasoning chains. In contrast, our DRP method remains effective even on such difficult tasks.
Existing SFT Methods struggle to balance accuracy and efficiency.
CoT-Valve achieves consistent token reductions across benchmarks by training on compressed CoT. However, this often comes with accuracy trade-offs. For instance, on MATH500 and AMC, we observe accuracy degradation despite reduced token usage. By comparison, our DRP method achieves both stronger compression and accuracy improvements across most benchmarks. Notably, on AMC (1.5B), DRP improves accuracy from 22/40 to 27/40 while also reducing tokens from 4574 to 3657.
Fine-grained supervised pruning improves reasoning accuracy more effectively than RL-based methods.
ThinkPrune demonstrates strong performance on the 1.5B model, suggesting that pruning is effective in eliminating distracting or redundant reasoning steps, particularly for smaller-capacity models. Notably, it achieves solid gains on benchmarks like AIME24 and AMC. However, our DRP method yields higher accuracy improvements, likely due to its use of high-quality teacher-guided pruning. While ThinkPrune achieves slightly better compression on a few tasks—due to its explicit optimization for token length, DRP achieves notably higher accuracy across the board, while still maintaining competitive compression rates.
5.3 Token Usage Distribution Shift
Figure 3: Normalized token length distributions across GSM8K, MATH500, AIME24, and AMC before and after SFT using the DeepSeek-R1-Distill-Qwen-7B model. The horizontal axis indicates the normalized token length (token count divided by the maximum allowed length), and the vertical axis represents the probability density. Blue curves correspond to our method (DRP), and orange curves denote the baseline. The reduction in long-tail completions and high-token outliers indicates that DRP mitigates verbose and degenerate reasoning failures, resulting in more robust and efficient inference.
Figure 3 shows the distribution of normalized token lengths before and after applying DRP across all benchmarks using R1-Distill-Qwen-7B. Detailed observations are discussed below:
Concise Reasoning Across Tasks.
DRP effectively compresses the overall reasoning length across all benchmarks. The main density of the token distribution shifts leftward, especially on harder out-of-domain tasks like MATH500, AIME24, and AMC. This demonstrates that DRP encourages more concise reasoning behavior.
Eliminating Verbose and Degenerate Reasoning Failures.
DRP significantly reduces long-tail completions that typically result from verbose or degenerate reasoning. On AMC, the baseline exhibits a bimodal distribution, where the secondary peak reflects unnecessarily lengthy yet still valid reasoning paths. More critically, we observe a clear drop in density near the upper token limit—especially in AIME and AMC—indicating that DRP mitigates degenerate cases where the model previously generated excessively long or looping outputs due to failure to converge (Section 4.6). This results in improved robustness, defined here as the model’s ability to terminate reasonably when it cannot reach a correct solution.
Further Compression of Already Efficient Reasoning Paths.
On datasets such as GSM8K and MATH500, where baseline models already produce relatively short completions, DRP still yields measurable compression gains. This indicates that DRP not only removes verbosity but also optimizes reasoning even in high-performing regimes.
6 Ablation Studies
We conduct three ablation experiments to understand the contribution of each component in our DRP framework separately.
6.1 Skill-based Decomposition vs. Default Step Split
RQ1: Does skill-based decomposition improve downstream learning compared to default step segmentation?
Dataset Method Accuracy Tokens
GSM8K 7B Base 91.7% 917
Default 92.7% 350
DRP 94.1% 328
OOD Tasks
MATH500 7B Base 92.4% 2486
Default 92.0% 1905
DRP 93.0% 1781
AIME24 7B Base 15/30 8674
Default 14/30 4678
DRP 15/30 4966
AMC 7B Base 31/40 4845
Default 31/40 4975
DRP 33/40 3258
Table 2: Comparison of 7B base model, default step segmentation, and our DRP method across in-distribution and out-of-distribution math benchmarks.
To assess the impact of our skill-based decomposition on downstream performance, we compare to a baseline that applies a default step-wise split without leveraging any functional skill information. Both variants are followed by teacher-guided revision and used to train models within the same DRP framework.
Comparison of Decomposition Prompts Skill-Based Prompt: Segment the explanation into a sequence of clear, non-overlapping steps, where each step corresponds to exactly one atomic mathematical skill. These skills may include, for example, addition, subtraction, applying a formula, interpreting a quantity, simplifying, checking a condition, and so on. Default Prompt: Segment the explanation into clear, non-overlapping steps.
As shown in Table 2, while default segmentation already provides some benefit under the self-revision setup, our skill-based segmentation consistently achieves stronger results across all benchmarks. DRP with skill-aware segmentation not only improves accuracy (e.g., +2 points on AMC) but also reduces token usage more effectively (e.g., 3258 vs. 4975 tokens on AMC).
This performance gap can be attributed in part to the difference in decomposition granularity: on the GSM8K training set, our method produces an average of 12.6 steps per example, compared to only 8.3 steps under the default strategy (see Appendix E for an illustrative example). The finer-grained, skill-aligned structure facilitates more precise pruning and better alignment with distinct math operations.
Crucially, this structural advantage also translates into improved generalization: we observe that the benefit of skill-based segmentation holds even on out-of-distribution reasoning tasks, further demonstrating its role in producing stable and transferable reasoning supervision.
Method GSM8K MATH500 AIME24 AMC
Accuracy Tokens Accuracy Tokens Accuracy Tokens Accuracy Tokens
7B-base student model 91.7% 917 92.4% 2486 15/30 8674 31/40 4845
   with GPT-4o teacher 94.1% 328 93.0% 1781 15/30 4966 33/40 3258
   with Gemini 2.0 flash teacher 93.2% 419 91.8% 2245 14/30 4860 33/40 2835
   with ChatGPT teacher 91.2% 386 90.2% 2122 14/30 5265 32/40 3101
Table 3: Impact of different teacher models on DRP performance. We compare GPT-4o, Gemini 2.0 Flash, and ChatGPT as pruning teachers, evaluating downstream accuracy and average token usage on four math benchmarks.
6.2 Structured Pruning vs. Direct Distillation
RQ2: Does the performance gain come from shorter CoTs or the structured pruning process?
Dataset Method Accuracy Tokens
GSM8K 7B Base 91.7% 917
Distill 90.7% 425
DRP 94.1% 328
OOD Tasks
MATH500 7B Base 92.4% 2486
Distill 88.6% 2152
DRP 93.0% 1781
AIME24 7B Base 15/30 8674
Distill 13/30 6417
DRP 15/30 4966
AMC 7B Base 31/40 4845
Distill 28/40 4279
DRP 33/40 3258
Table 4: Comparison between 7B base model, direct distillation from GPT-4o, and our DRP method across both in-distribution and OOD math benchmarks.
To answer this, we compare DRP with a direct distillation baseline, where the student model is trained on concise CoT generated directly by GPT-4o. These responses are notably short—averaging 186 tokens—compared to DRP’s CoT average of 330 tokens. Table 4 presents the results.
We observe that:
  • On the in-distribution dataset GSM8K, direct distillation reduces token usage by over 50% (425 vs. 917 tokens) with only a minor accuracy drop (90.7% vs. 91.7%), indicating its effectiveness in simpler tasks.
  • On OOD tasks like MATH500, AIME24, and AMC, however, accuracy drops significantly despite only modest token savings (e.g., 88.6% vs. 93.0% on MATH500; 13/30 vs. 15/30 on AIME24), suggesting limited generalization.
These findings suggest that shorter CoT alone do not guarantee better generalization in reasoning tasks. While direct distillation from teacher LLMs provides compressed outputs, these often omit essential intermediate reasoning steps. In contrast, DRP explicitly removes redundant content while preserving a step-level, skill-aligned structure, allowing the model to internalize more transferable problem-solving patterns.
This highlights an important insight: Retaining structurally meaningful reasoning—even at the cost of moderate length—yields superior downstream performance compared to blindly imitating shorter outputs. A comparison between short-form and DRP-style trajectories is in Appendix F.
6.3 Influence of Teacher Model Choice
RQ3: How does the choice of teacher model influence pruning effectiveness?
To evaluate the sensitivity of our pruning framework to the choice of teacher model, we experiment with three different large language models (LLMs): GPT-4o, Gemini 2.0 Flash, and ChatGPT. As shown in Table 3, all teacher models lead to consistent improvements in both accuracy and token efficiency compared to the base model, confirming the robustness and generalizability of our approach.
Among them, GPT-4o yields the strongest performance across most benchmarks, particularly in token compression. However, the overall differences between teacher models remain modest, suggesting that while stronger teachers can produce slightly more optimal pruning decisions, our method remains effective even when relying on weaker or more accessible models.
This indicates that our framework does not overfit to a specific teacher’s internal reasoning patterns, and can generalize pruning supervision across a variety of LLMs. Notably, we observe that variations across teachers are more pronounced in token usage than in accuracy—highlighting that our method’s core benefit, skill-level reasoning distillation, remains stable regardless of teacher strength.
7 Conclusion
We propose skill-based Distilled Reasoning Pruning, a framework that leverages pruned reasoning traces to distill smaller reasoning models. It outperforms direct distillation by effectively bridging the learnability gap between student and teacher models. The success of our skill-based step decomposition underscores the importance of fine-grained, consistent step segmentation as a strong foundation for pruning.
8 Limitations
While our method demonstrates strong performance on existing small-scale reasoning models, it remains uncertain how well it generalizes to other architectures. Currently, there is a limited number of publicly available small-sized LRMs with strong reasoning capabilities, making broad validation challenging. In addition, the training paradigms for reasoning models are evolving rapidly, and it is unclear whether the overthinking and inefficiency issues we target will persist in future model generations. We view our work as a step toward addressing current bottlenecks, but acknowledge that its relevance may shift as the landscape of LLM training continues to change.
9 Ethics
Our study uses the OpenAI and Google Gemini API on experiments above, and at no point do we access, or attempt to access, the true training data behind these models, or any underlying components of the systems. Risks The several datasets in our experiment are sourced from publicly available sources. However, we cannot guarantee that they are devoid of socially harmful or toxic language. We use ChatGPT  [^4] https://chatgpt.com/ to correct grammatical errors in this paper.
Appendix A Outliers and Cutting Off
We can observe from Figure 4 that most answers are captured by 11K tokens, though some require up to 38K, revealing a long-tail pattern.
Figure 4: The x-axis denotes the model’s maximum generation length, and the y-axis shows the proportion of correct answers recovered within that budget, with the R1-Distill-Qwen-7B.
Method GSM8K MATH500 AIME24 AMC
Accuracy Tokens Accuracy Tokens Accuracy Tokens Accuracy Tokens
Baseline (no cutoff) 91.7% 917 92.5% 5435 15/30 54339 32/40 18736
Baseline (12k cutoff) 91.7% 917 92.4% 2486 15/30 8674 31/40 4845
Table 5: Impact of applying a 12k token cutoff on measured token usage across benchmarks. Accuracy remains largely unaffected, while the average token count drops significantly—particularly on harder tasks like AIME and AMC. This demonstrates the necessity of outlier mitigation for fair efficiency comparison.
As shown in Table 5, applying the 12k token cutoff has a significant effect, particularly on harder benchmarks like AIME24 and AMC. On these tasks, the model occasionally fails to generate valid answers and enters degenerate loops, repeatedly producing the same text until reaching the maximum token limit. As a result, the average token usage can increase by up to 5×, which does not reflect the true distribution of reasoning length and severely skews efficiency comparisons.
Appendix B Fine-Tuning Details
We fine-tune the DeepSeek-R1-Distill-Qwen models (7B and 1.5B) using the LLaMA-Factory framework with LoRA adaptation. The training data includes 8,000 samples drawn from GSM8K and PRM12K. We set a cutoff length of 4096 tokens for both input and output sequences.
The models are fine-tuned for 3 epochs using the following configuration:
  • Cutoff length: 4096
  • Max samples: 8000
  • Batch size: 2 (with gradient accumulation of 4)
  • Learning rate: 3e-5 with cosine schedule
  • Precision: bf16
  • Validation split: 5% of training data
  • Evaluation strategy: every 300 steps
Training is performed using two A100 80GB GPU. All experiments use overwrite_cache=true and 8 parallel preprocessing workers. The resulting models are directly used for downstream evaluation without additional tuning.
Appendix C Prompt Templates
The full prompt for skill-based step segmentation:
Skill-Based Step Segmentation Prompt (Full Version) You are given a complete reasoning path for a math problem. Your task is to segment it into a sequence of clear, non-overlapping steps, where each step corresponds to exactly one atomic mathematical skill. These skills may include, for example, addition, subtraction, applying a formula, interpreting a quantity, simplifying, checking a condition, and so on. Use the following format for each step: Step n: {{original text segment}}
Skill: {{name of the skill used}}
Only segment and label the steps — do not solve or modify the original content in any way.
The full prompt for self-revision:
Self-Revision Prompt (Full Version) You are an expert in mathematical reasoning compression. Given a list of reasoning steps labeled with their respective skills, your task is to evaluate and revise each step according to one of the following four actions: 1. KEEP: The step is necessary and already concise. Keep it unchanged.
2. DELETE: The step is unnecessary and should be removed entirely.
3. SINGLE-STEP COMPRESS: The step is necessary but verbose; rewrite it in a more concise way.
4. MULTI-STEP COMPRESS: The step can be merged with neighboring steps; write a combined, cleaner version.
If the final step clarifies the final answer (e.g., “The answer is…”), retain it. After completing the revision of each step, synthesize the revised steps into a coherent explanation. Ensure the output is fluent, logically sound, and matches the speaker’s tone and style.
Appendix D Skill Base Decomposition Evaluation
To assess the quality of our skill-based step decomposition, we conduct a pairwise comparison experiment using Gemini 2.0 Flash as the judge. For each reasoning trace $T$, we generate two segmentations: (1) our proposed skill-based decomposition, and (2) a baseline sentence-based split (obtained by naïvely splitting at punctuation or sentence boundaries).
We prompt Gemini to select the version with more semantically coherent, structurally stable, and granularity-appropriate steps, using the following instruction:
Given two versions of step-by-step reasoning derived from the same original trace, please choose the version that shows better semantic coherence, structural consistency, and appropriate granularity. Respond with ‘A’, ‘B’, or ‘Same’.
Among 50 randomly sampled traces from GSM8K, Gemini selected our skill-based decomposition in 33 cases, while preferring the baseline in only 17 cases. This result suggests that our segmentation method provides more interpretable and structurally aligned reasoning steps, which are beneficial for downstream pruning and distillation.
Appendix E Skill-based Decomposition vs. Default
In this section, we show a non-cherry picking example of how our skill based decomposition different from the default one. As shown in Table 6, our method has a more fine-grained step decomposition effect which leads to a more detailed pruning effect.
Form Decomposition Output
Problem Joy can read 8 pages of a book in 20 minutes. How many hours will it take her to read 120 pages?
Original Long-form CoT First, I need to determine Joy’s reading rate. She reads 8 pages in 20 minutes, so her reading rate is 8 pages per 20 minutes. Next, I’ll calculate how many pages she reads per minute by dividing 8 pages by 20 minutes, which gives 0.4 pages per minute. Then, to find out how many minutes it takes her to read 120 pages, I’ll divide 120 pages by her reading rate of 0.4 pages per minute, resulting in 300 minutes. Finally, I’ll convert 300 minutes into hours by dividing by 60, which equals 5 hours.
DRP Decomposition (Skill-Based) Step 1: Determine Joy’s reading rate. She reads 8 pages in 20 minutes, so her reading rate is 8 pages per 20 minutes.
Skill: Interpreting a quantity
Step 2: Calculate how many pages she reads per minute by dividing 8 pages by 20 minutes.
Skill: Division
Step 3: The result of the division is 0.4 pages per minute.
Skill: Simplifying
Step 4: To find out how many minutes it takes her to read 120 pages, divide 120 pages by 0.4 pages per minute.
Skill: Division
Step 5: The result of the division is 300 minutes.
Skill: Simplifying
Step 6: Convert 300 minutes into hours by dividing by 60.
Skill: Division
Step 7: The result of the division is 5 hours.
Skill: Simplifying
DRP Pruning and Final Result 1. KEEP: The step is necessary and already concise.
2. SINGLE-STEP COMPRESS: Joy reads 8 pages in 20 minutes, so she reads 0.4 pages per minute (8 ÷ 20).
3. DELETE: This step is now included in the previous step.
4. SINGLE-STEP COMPRESS: To find the time for 120 pages, divide 120 by 0.4 pages per minute.
5. DELETE: This step is now included in the previous step.
6. SINGLE-STEP COMPRESS: Convert 300 minutes to hours by dividing by 60.
7. DELETE: This step is now included in the previous step.
Final Answer: Okay, let’s calculate the time it takes Joy to read 120 pages. Joy reads 8 pages in 20 minutes, which means she reads 0.4 pages per minute (8 ÷ 20). To find out how many minutes it takes her to read 120 pages, divide 120 by 0.4, resulting in 300 minutes. Convert 300 minutes into hours by dividing by 60, which gives us 5 hours. The answer is 5 hours.
Default Decomposition Step 1: First, I need to determine Joy’s reading rate. She reads 8 pages in 20 minutes, so her reading rate is 8 pages per 20 minutes.
Step 2: Next, I’ll calculate how many pages she reads per minute by dividing 8 pages by 20 minutes, which gives 0.4 pages per minute.
Step 3: Then, to find out how many minutes it takes her to read 120 pages, I’ll divide 120 pages by her reading rate of 0.4 pages per minute, resulting in 300 minutes.
Step 4: Finally, I’ll convert 300 minutes into hours by dividing by 60, which equals 5 hours.
Default Pruning and Final Result Step 1: SINGLE-STEP COMPRESS
Revised: Joy’s reading rate is 8 pages per 20 minutes.
Step 2: SINGLE-STEP COMPRESS
Revised: Dividing 8 by 20, Joy reads 0.4 pages per minute.
Step 3: SINGLE-STEP COMPRESS
Revised: Dividing 120 by 0.4, it takes her 300 minutes to read 120 pages.
Step 4: SINGLE-STEP COMPRESS
Revised: Converting 300 minutes to hours: 300 ÷ 60 = 5 hours.
Final Answer:
Joy’s reading rate is 8 pages per 20 minutes. Dividing 8 by 20, Joy reads 0.4 pages per minute. Dividing 120 by 0.4, it takes her 300 minutes to read 120 pages. Converting 300 minutes to hours: 300 ÷ 60 = 5 hours. The answer is 5.
Table 6: A comparison between DRP’s skill-based decomposition and default segmentation with compression, for the same math problem.
Appendix F Long CoT vs. Short CoT
In this section, we show a non-cherry picking example of how our skill based decomposition different from the default one. As shown in Table 7, our DRP-style CoT are more informative. That might because of LLM’s unfair ability on different aspects Jiang and Ferraro (2024).
Form Reasoning Trajectory
Problem: Joy can read 8 pages of a book in 20 minutes. How many hours will it take her to read 120 pages?
Original Long-form CoT First, I need to determine Joy’s reading rate. She reads 8 pages in 20 minutes, so her reading rate is 8 pages per 20 minutes. Next, I’ll calculate how many pages she reads per minute by dividing 8 pages by 20 minutes, which gives 0.4 pages per minute. Then, to find out how many minutes it takes her to read 120 pages, I’ll divide 120 pages by her reading rate of 0.4 pages per minute, resulting in 300 minutes. Finally, I’ll convert 300 minutes into hours by dividing by 60, which equals 5 hours.
DRP-style (Ours) Okay, let’s calculate the time it takes Joy to read 120 pages. Joy reads 8 pages in 20 minutes, which means she reads 0.4 pages per minute ($8\div 20$). To find out how many minutes it takes her to read 120 pages, divide 120 by 0.4, resulting in 300 minutes. Convert 300 minutes into hours by dividing by 60, which gives us 5 hours. The answer is 5 hours.
Short-form CoT In one hour, there are 3 sets of 20 minutes. So, Joy can read $8\times 3=24$ pages in an hour. It will take her $120\div 24=5$ hours to read 120 pages.
Table 7: A comparison between original long-form, DRP-style, and short-form reasoning traces for the same math problem.